OpenStudy (idealist10):

Simplify: u"e^(x)-2u"xe^(x)-4u'xe^(x)+4u'e^(x)=(1-4x+4x^2)e^x

OpenStudy (idealist10):

\[u''e^x-2u''xe^x-4u'xe^x+4u'e^x=(1-4x+4x^2)e^x \]

OpenStudy (idealist10):

@freckles @thomaster

OpenStudy (pawanyadav):

u"e^x(1-2x) +4u'e^x(1-x)=(1-4x-4x^2)e^x Divide both side by e^x =u"(1-2x) +u'(1-x)= (1-4x-4x^2)

OpenStudy (pawanyadav):

Now multiply both side by dx then integrate it

OpenStudy (idealist10):

@mathstudent55